27cm2 : 57cm2 = 18cm : 38cm
 27cm2 : 57cm2 = 18cm : 38cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.
Consider the two ratios 27cm^2: 57cm^2= 18cm : 38cm

27/57 = 18/38
27/57 is further simplified by dividing both numerator and denominator by 3
we get, = 9/19
18/38 is further simplified by dividing both numerator and denominator by 2
we get, = 9/19
Hence, 9/19 = 9/19
Therefore, 27cm^2: 57cm^2= 18cm : 38cm

Thanks